Consider the following hypotheses:
H0: p ≥ 0.47
HA: p < 0.47

Compute the p-value based on the following sample information. (You may find it useful to reference the appropriate table: z table or t table) (Round "z" value to 2 decimal places. Round intermediate calculations to at least 4 decimal places and final answers to 4 decimal places.)
p value

Consider the following hypotheses:
H0: p ≥ 0.47
HA: p < 0.47

Compute the p-value based on the following sample information. (You may find it useful to reference the appropriate table: z table or t table) (Round "z" value to 2 decimal places. Round intermediate calculations to at least 4 decimal places and final answers to 4 decimal places.)

p value
x = 112; n = 311 _______

Answer by Boreal(15235)

112/311=0.360
z=(phat-0.47)/sqrt(0.47*0.53/311); the sqrt term is the std error.
=-0.11/0.0283
=-3.89
p-value is the amount -3.89 is from the left side of the curve (infinite left) or prob z < -3.89
it is 0.000052
p<0.0001 would normally be sufficient but use the above as the answer.